Last week (Friday 3 February) the LCO education team hosted a teacher training event in Cardiff University, jointly with School of Physics & Astronomy, Universe in the Classroom, and Space Awareness. It was the not only the largest teacher training event ever run at the School of Physics & Astronomy (with 35 registered teachers from across Wales) but was also our first event as official partners with Space Awareness.
Following the recommendations of an independent Time Allocation Committee (TAC), five new Key Projects have been awarded time on the LCO Network. The list of all key projects that will run during the 2017AB semester has been posted on the website's Science section.
Las Cumbres Observatory is soliciting proposals for science observations for the 2017AB semester, which will begin on 1 April 2017 and run through 30 November 2017. This call is for astronomers from institutions without guarantees of Network time: the University of California at Santa Barbara (UCSB), the Infrared Processing and Analysis Center (IPAC), the University of Colorado, and the Instituto de Astrofisica de Canarias (IAC). Astronomers affiliated with LCO, including Time Allocation Committee members and Science Advisory Committee members, are also welcome to submit proposals. (LCO's Science Collaboration partners that have guaranteed time conduct independent proposal solicitations and reviews.)
The Network suffered a one-two punch of power outages in the past week. On December 17, power was lost on Haleakala. As a result, science observations from that site (OGG) weren't possible on December 17-18. ----- A few days earlier, on December 13, there were power outages that affected all CTIO facilities. Consequently, we made no science observation from our LSC site on December 13-14. Since then, the cameras in Dome B (fl03) and Dome C (fl04) at LSC have been showing unstable temperatures. As a safeguard, we will warm these cameras up tonight (2016-12-21), then inspect and recharge their cryogenic systems tomorrow. Thus, only one camera (fl15, in Dome A) will be available for science observations tonight. We expect to have all cameras available again for the night of 2016-12-22.
A version of this article is also available for young scientists! Read it at
As well as being the LCO education team, Sarah Eve Roberts and I are also the team behind Space Scoop, an astronomy news service for kids. Although the primary audience for Space Scoop is kids from the age of 8, we know that many older kids and adults also enjoy reading Space Scoop. Each Scoop is written about a brand discovery in astronomy or space sciences, based on a current press release. We aim to release every Space Scoop at the same time as the research press release is published, that way scientists, journalists and kids all get to hear about exciting space news at the same time.
A new SBIG camera (kb98) was installed on the 0m4a (40cm) telescope at our Australia site on November 24. The camera was enabled for science observations on December 1.
Also at the Australia site, the 2-meter telescope was taken offline on December 4. Our TelOps team detected some anomalous behavior in the telescope's motion. They've now found the likely source of the anomaly and fixed the problem. We expect to re-enable the telescope for science observation for the 2016-12-08 night.
I reported in November that we've been experiencing internet connectivity problems to our site in Chile. Those problems, although now less frequent, haven't ceased. The unreliable connection can disrupt the delivery (and subsequent processing) of images. If you can't find images in the archive from observations made at our Chile site (LSC), please send an email to email@example.com
In the past few weeks, we've experienced intermittent problems with internet connectivity to some of our network sites, which has prevented us from promptly delivering images to our science archive. In all of these cases, the connectivity problems were not due to LCO hardware or software failures, but to failures beyond our power to fix (e.g. fiber breaks). Connections to Australia and South Africa have now been stable for several days, but our connection to Chile remains unreliable. The connectivity problems are affecting all facilities on Cerro Tololo.
One week ago, all of LCO's 1-meter telescopes are equipped with Sinistro cameras. Today, all of the Sinistros are available for science observations, and the raw images are getting processed through the BANZAI reduction pipeline.
The shutter on the Spectral camera on FTN (Haleakala) was replaced last week, and test observations indicate that the shutter is behaving itself. The camera has been cleared to resume science observations.
Subscribe to our news RSS /